Pitney Co. purchased an office building, land, and furniture for $692,200 cash. The appraised value of the assets was as follows: Land $ 131,795 Building 240,332 Furniture 403,137 Total $ 775,264

Required

  1. Compute the amount to be recorded on the books for each asset.

  1. Record the purchase in a horizontal statements model.
Answer #1
a
Allocated cost
Land 117674 =692200*131795/775264
Building 214582 =692200*240332/775264
Furniture 359944 =692200*403137/775264
Total 692200
b
Assets Equity Revenue Expense Net income Cash flows
Cash Land Building Furniture
-692200 117674 214582 359944 -692200 IA